OpenStudy (anonymous):

The area of a rectangular garden is given by the trinomial x^2 + 6x – 27. What are the possible dimensions of the rectangle? Use factoring.

OpenStudy (bahrom7893):

x^2+6x=27 x(x+6)=27 x=3 or x=9

OpenStudy (bahrom7893):

wait ignore the last line

OpenStudy (bahrom7893):

x^2+6x=27 x(x+6)=27 x = 3 so one side could be 3, the other could be 9.

OpenStudy (anonymous):

x^2+6x+27=0 x^2+3x+9x+27=0 x(x+3)+9(x+3)=0 x+3=0 or x+9=0 x=-3 x=-9

OpenStudy (bahrom7893):

it was -27 StAr

OpenStudy (anonymous):

Woops sorry !

OpenStudy (anonymous):

x^2+6x-27=0 x^2-3x+9x-27=0 x(x-3)+9(x-3)=0 x-3=0 or x+9=0 x=3 x=-9

OpenStudy (anonymous):

thanks you guys.